OpenStudy (anonymous):

x=y^2-1 2x-y=4 Which of the following equations could be the result of using substitution to solve the system? 2y^2-y-5=0 2y^2-y-6=0 2y^2-2y-5=0

OpenStudy (anonymous):

solve 2x-y=4 for x and plug it into x=y^2-1

OpenStudy (anonymous):

2y^2-y-6=0

OpenStudy (anonymous):

Thanks :D

OpenStudy (anonymous):

np